OpenStudy (nali):

find the standard form of the equation of the parabola with vertex at the origin focus (-2,0)

OpenStudy (nali):

plz someone help me in getting started and i can do the rest

OpenStudy (anonymous):

all parabolas with its vertex at the origin has this form: 4py = x^2 up/down parabola 4px = y^2 left/right parabola where p is the distance from vertex to focus. if p is negative, it opens downard or left.

OpenStudy (anonymous):

so with your focus being at (-2, 0), this is a parabola that opens to the left... correct?

OpenStudy (anonymous):



OpenStudy (anonymous):

so it should have equation: 4px = y^2 since the distance from the vertex to the parabola is 2, and it opens left, then 4(-2)x = y^2 or -8x = y^2 that's the equation of the parabola whose vertex is at the origin and focus at (-2, 0).

OpenStudy (nali):

thanku so much